# Large Language Models for Scientific Idea Generation: A Creativity-Centered Survey

### v1

@@ 7 Evaluation of Scientific Idea Generation: Frameworks and Challenges
Evaluating LLM-generated scientific ideas is crucial yet difficult. Although evaluation mechanisms share techniques and sources with the feedback systems discussed in Section 4.2, they serve different purposes: feedback guides idea generation through search and refinement, while evaluation assesses final outputs to determine a model’s effectiveness as a scientific ideator. In Rhodes’ 4Ps (rhodes1961analysis) framework, the Product dimension—novel, useful, and correct outputs—serves as the key endpoint of creativity, but assessing it is challenging because scientific ideas are open-ended and their true impact unfolds over time. This makes evaluation a major bottleneck: expert judgment is insightful but unscalable, automated metrics are efficient but reductive, and LLM-as-a-judge approaches promise a balance while still inheriting biases. In this section, we survey three major families of evaluation frameworks: (1) Computational and Execution-based metrics, which operationalize dimensions of creativity such as novelty, diversity, feasibility and utility into quantifiable scores; (2) Expert review, which captures nuanced human intuition beyond what metrics alone can measure; and (3) LLM-as-a-judge models, which aim to synthesize the advantages of automation and human judgment. Table 7.1 summarizes Evaluation paradigms, highlighting their core strategies, strengths, and limitations.
@@ 7.1 Computational and Execution-Based Metrics
Computational and Execution-Based Metrics translate creativity dimensions—novelty, diversity, feasibility, and impact—into measurable, reproducible signals. Computational metrics derive these signals analytically through Model-based formulations, while execution-based metrics validate them empirically through Simulation or real-world testing. Early work, such as SciMon (wang2024scimon), evaluates the overall quality of generated ideas using surface-level text similarity measures (e.g., rouge, bertscore, bartscore), without distinguishing between different aspects of creativity to compare generated ideas against reference texts. While such holistic NLP-based evaluations offer scalability and ease of automation, they fail to capture the conceptual depth and open-ended multidimensional nature of scientific creativity.
@@
Table 3: Comparison of evaluation paradigms for scientific idea generation. The table summarizes representative methods, example works, their strengths and limitations, and the Best captured Creativity Dimension (BCCD).
[…]
Human Judgment • Expert panel, real peer-review submission • Consensual Assessment Technique • Structured Likert-scale
Scideator(smith2025scideator**), Nova(hu2024nova),
AI Scientist v2(yamada2025aiscientistv2)
CAT(Amabile1982social)
AIdeation10.1145/3706598.3714148
+ Conceptual depth, captures nuanced aspects - Time-consuming, costly, subjective, un-scalable Feasibility, Impact
[…]
• Prompt-based or inference-time judges
• Fine-tuned judges trained on peer-review data
• Reinforcement-optimized reward models
• Hybrid frameworks
LiveIdeaBench (ruan2025liveideabenchevaluatingllmsdivergent),
AI Scientist (lu2024aiscientistfullyautomated)
CycleReviewer (weng2025cycleresearcher),
OpenReviewer idahl-ahmadi-2025-openreviewer
HARPA (vasu2025harpatestabilitydrivenliteraturegroundedframework),
ReviewRL(zeng2025reviewrl)
DeepReviewer(zhu2025deepreview)
+ Scalable, reproducible, approximates expert reasoning and intuition
- Sensitive to prompts, training bias, and domain drift
Feasibility, Novelty, Impact
[…]
• Text similarity metrics: BLEU, ROUGE, BERTScore, DSI
• Semantic distance RND, HD, ON, APD, Vendi Score
• Graph-based, Network Centrality • Retrieval or Bayesian fusion
SciMon (wang2024scimon), CrPO(2025crpo)
VirSci(su2024virsci),
RND(wang2025enablingaiscientistsrecognize),
Vendi Scorefriedman2023vendiscorediversityevaluation
SciND(gupta2024scind) (ke2023networknormalizedimpact)
HypoAgent(duan2025bayesentropycollaborativedrivenagents)
+ Objective, interpretable, domain-agnostic, scalable
- Surface-level semantics; depends on corpus quality and embedding fidelity
Novelty, Diversity, Impact
[…]
• Simulators, Compilers, Autonomous Labs
• Behavioral execution
AI Scientist (lu2024aiscientistfullyautomated),
AI Co-Scientist(gottweis2025aicoscientist)
MindShift (Wu_2024),
LLM Counter Speech(podolak2024llmgeneratedresponsesmitigate)
+ Grounded, outcome-based validation
- Expensive, domain-limited, low scalability
Feasibility, Impact
[…]
@@ 7.1.1 Novelty and Originality
Several recent studies have approached novelty assessment through semantic local density, assuming that truly novel ideas should lie in sparse regions of the scientific embedding space. su2024virsci introduced two such measures: Historical Dissimilarity (HD)—the average Euclidean distance between a generated idea’s embedding and its five most similar historical abstracts—and Overall Novelty (ON), which further incorporates the contemporaneity and impact of neighboring works:
@@
Building on this intuition, Relative Neighbor Density (RND) evaluates novelty through the density of an idea’s semantic neighborhood relative to its neighbors’ own densities (wang2025enablingaiscientistsrecognize). Given a set of ideas $I={{\text{idea}_{i}}}$, each represented by an embedding $v_{i}=G(\text{idea}_{i})$ obtained via a semantic encoder $G(\cdot)$, RND computes for each idea $i$ its $P$ nearest neighbors in the literature corpus $A$:
@@ 7.1.2 Diversity
While novelty measures how far a single idea departs from prior knowledge, diversity evaluates the breadth of a set of generated ideas. A model that outputs only a narrow cluster of novel but similar hypotheses may underexplore the solution space. Diversity thus ensures that systems span multiple conceptual directions, not just isolated rare points. Empirical studies confirm that high novelty does not imply diversity—models may generate unique yet domain-concentrated ideas—underscoring the need to evaluate both dimensions jointly (ruan2025liveideabenchevaluatingllmsdivergent).
@@
Building on the same principle, instead of evaluating the whole set at once, CrPO 2025crpo, computes a per-response dispersion score. Responses with higher score—i.e., those further from their peers—receive stronger rewards, encouraging models to cover multiple directions rather than collapsing to a single theme.
Beyond cosine-based measures, Vendi Score friedman2023vendiscorediversityevaluation provides a more principled measure of set distinctness using Shannon entropy over the eigenvalues of the similarity matrix. Given a similarity matrix $K\in\mathbb{R}^{M\times M}$ constructed from embeddings, let $\{\lambda_{1},\dots,\lambda_{M}\}$ denote its normalized eigenvalues (summing to 1). The Vendi Score is defined as:
@@ 7.1.3 Feasibility and Resource-Efficiency
The HypoAgents framework (duan2025bayesentropycollaborativedrivenagents) proposes an N–R–F (Novelty–Relevance–Feasibility) model that quantifies idea quality via Bayesian inference and information fusion. Using retrieved literature and semantic embeddings, it measures feasibility as the Bayesian probability of supporting evidence—deriving scores computationally.
Execution-grounded evaluations appear most clearly in autonomous research pipelines. For instance, AI Scientist papers (lu2024aiscientistfullyautomated; yamada2025aiscientistv2) or DeepScientist (weng2025deepscientist) integrates idea generation, code synthesis, and experimental validation in a closed loop. Hypotheses proposed by one agent are automatically implemented as machine-learning experiments, trained, and benchmarked; success metrics such as performance gain or reproducibility indicate feasibility. Similarly, HARPA (vasu2025harpatestabilitydrivenliteraturegroundedframework) introduces a testability-driven ideation framework, where an LLM reward model predicts feasibility scores and selected hypotheses are subsequently verified through simulation or real-world trials, closing the loop between idea generation and empirical testing.
In materials and molecular science, execution-based feasibility is operationalized through simulation and experimental pipelines. borg2023quantifying introduce empirical measures such as Discovery Yield—the fraction of target materials discovered—and Discovery Probability—the likelihood of identifying at least one viable candidate—both derived from iterative experimental cycles. Similarly, gallagher2024data quantify efficiency as the proportion of candidate compounds passing synthesizability, stability, or toxicity constraints during virtual screening, effectively measuring the fraction of computationally or experimentally “non-wasted” recommendations. Systems like BenchML (poelking2021benchmlextensiblepipeliningframework) combine simulators and synthesis predictors to test chemical hypotheses, evaluating feasibility based on success rates such as molecular stability or docking performance.
@@ 7.1.4 Utility and Impact
where $\sigma_{st}$ is the total number of shortest paths between nodes $s$ and $t$, and $\sigma_{st}(y)$ is the number of such paths passing through $y$. High $C_{B}(y)$ indicates that the idea bridges otherwise disconnected communities, a property correlated with transformative scientific discoveries (ke2023networknormalizedimpact). Such metrics provide a scalable, domain-agnostic way to estimate potential influence before real deployment, though they cannot capture context-specific or practical impact.
@@
The AI Co-Scientist (gottweis2025aicoscientist) explicitly executes biomedical hypotheses to quantify real impact through end-to-end wet-lab experimentation. It autonomously proposes and tests hypotheses in three increasingly complex domains—drug repurposing, novel target discovery, and microbial evolution. The system systematically measures downstream biological outcomes—such as inhibition efficacy or regenerative response—to assess the practical significance of its discoveries. This rigorous execution-driven evaluation exemplifies how impact in scientific ideation can be assessed through direct experimentation rather than proxy metrics.
Beyond laboratory science, behavioral execution studies such as MindShift (Wu_2024) and LLM Counter-Speech (podolak2024llmgeneratedresponsesmitigate) deploy AI-generated interventions in natural environments—digital health and social media—to quantify real behavioral change as a domain-specific key performance indicator (KPI) measuring practical impact. Similarly, large-scale agent-based frameworks like AgentSociety (piao2025agentsocietylargescalesimulationllmdriven) model how new social or policy ideas propagate through simulated societies, tracking collective outcomes as impact proxies. In summary, impact assessment through execution-driven tests, behavioral outcome studies or graph-based scientometric, ensure that models are not merely producing ideas that are novel, diverse, or feasible, but ones that are also meaningful and capable of advancing science or society in a measurable way. However, such approaches are inherently limited: many scientific ideas are not directly executable, domain simulators exist only for select fields, and measurable outcomes often capture only the practical success of an idea rather than its deeper creative or theoretical significance. Consequently, nuanced dimensions of creativity and long-term scientific value still rely on human expertise and intuition, which remain the gold standard for assessing scientific ideas.
Table 7.1 summarizes the main evaluation paradigms—from human judgment to computational and execution-based methods—and their roles in assessing novelty, diversity, feasibility and impact.
@@ 7.2 Human Expert Evaluation
Typical setups follow a panel review format, modeled after peer review. For example, AIdeation (10.1145/3706598.3714148) engaged professional concept designers to rate generated ideas using Likert scales (e.g., 1–5 ratings) for creativity, feasibility, and satisfaction, supplemented by qualitative interviews. Similar setups are found in Scideator (smith2025scideator**) and Nova (hu2024nova), where domain specialists score hypotheses for novelty, plausibility, and usefulness, providing both quantitative and qualitative human feedback.
To improve inter-rater reliability, several studies adopt the Consensual Assessment Technique (CAT) (Amabile1982social), aggregating multiple expert scores into consensus ratings that reduce bias and align evaluation with established creativity research. At the most rigorous level, AI Scientist v2 (yamada2025aiscientistv2) elevates human evaluation to formal peer review by submitting autonomously generated papers to top-tier conferences, where acceptance serves as the strongest validation of impact and scientific merit.
@@
[…]
Inference-time judges (Prompt-based) Leverages structured prompts, retrieval engines or multi-agent critique to approximate expert review without fine-tuning.
AI Scientist (lu2024aiscientistfullyautomated)
LiveIdeaBench (ruan2025liveideabenchevaluatingllmsdivergent) Chain-of-Idea(li2024chain)
Low-cost, interpretable, and reproducible; easy to generalize across domains and tasks. Highly sensitive to prompt phrasing and context; lacks consistent calibration across runs.
Supervised Fine-Tuned Judges Trains models on annotated peer-review corpora to internalize expert-like judgment and scoring patterns.
CycleReviewer (weng2025cycleresearcher)
OpenReviewer (idahl-ahmadi-2025-openreviewer)
High correlation with expert judgments, Improved consistency and interpretability. Memorization instead of reasoning, Static preferences, Dependent on dataset coverage and quality; may inherit bias.
RL and Reward-Tuned Judges Optimizes evaluative behavior via reinforcement learning or preference alignment for reasoning stability.
HARPA (vasu2025harpatestabilitydrivenliteraturegroundedframework)
ReviewRL (zeng2025reviewrl)
REMOR (taechoyotin2025remorautomatedpeerreview)
Generalization and Adaptivity High correlation with expert judgments; encourages coherent, stable, and self-consistent evaluations. Expensive to train; vulnerable to reward mis-specification and preference drift.
Hybrid (Training + Inference Integration) Combines fine-tuning with inference-time reasoning and external verifiers or factuality tools.
DeepReview (zhu2025deepreview)
HARPA (vasu2025harpatestabilitydrivenliteraturegroundedframework)
Merge the generalization of training-based methods with the factual grounding of Inference scaling methods Complex pipeline; requires maintaining multi-stage logic and external dependencies.

### v2

@@ 7 Evaluation of Scientific Idea Generation: Frameworks and Challenges
Evaluating LLM-generated scientific ideas is crucial yet difficult. Although evaluation mechanisms share techniques and sources with the feedback systems discussed in Section 4.2, they serve different purposes: feedback guides idea generation through search and refinement, while evaluation assesses final outputs to determine a model’s effectiveness as a scientific ideator. In Rhodes’ 4Ps (Rhodes, 1961) framework, the Product dimension—novel, useful, and correct outputs—serves as the key endpoint of creativity, but assessing it is challenging because scientific ideas are open-ended and their true impact unfolds over time. This makes evaluation a major bottleneck: expert judgment is insightful but unscalable, automated metrics are efficient but reductive, and LLM-as-a-judge approaches promise a balance while still inheriting biases. In this section, we survey three major families of evaluation frameworks: (1) Computational and Execution-based metrics, which operationalize dimensions of creativity such as novelty, diversity, feasibility and utility into quantifiable scores; (2) Expert review, which captures nuanced human intuition beyond what metrics alone can measure; and (3) LLM-as-a-judge models, which aim to synthesize the advantages of automation and human judgment. Table 4 summarizes Evaluation paradigms, highlighting their core strategies, strengths, and limitations.
@@ 7.1 Computational and Execution-Based Metrics
Computational and Execution-Based Metrics translate creativity dimensions—novelty, valueness (diversity, feasibility, and impact)—into measurable, reproducible signals; however, we emphasize that these metrics provide indirect proxies for creativity rather than direct measurements of creative insight, and should be interpreted as operational indicators rather than exhaustive characterizations. Computational metrics derive these signals analytically through Model-based formulations, while execution-based metrics validate them empirically through Simulation or real-world testing. Early work, such as SciMon (Wang et al., 2024a), evaluates the overall quality of generated ideas using surface-level text similarity measures (e.g., rouge, bertscore, bartscore), without distinguishing between different aspects of creativity to compare generated ideas against reference texts. While such holistic NLP-based evaluations offer scalability and ease of automation, they fail to capture the conceptual depth and open-ended multidimensional nature of scientific creativity.
@@
Table 4: Comparison of evaluation paradigms for scientific idea generation.
Evaluation paradigm Representative methods Example works Strengths / limitations Creativity Dimension
Human Judgment • Expert panel, real peer-review submission • Consensual Assessment Technique • Structured Likert-scale Scideator (Smith and Huang, 2025), Nova (Hu et al., 2024), AI Scientist v2 (Yamada et al., 2025), CAT (Amabile, 1982), AIdeation (Wang et al., 2025d) + Conceptual depth, captures nuanced aspects - Time-consuming, costly, subjective, unscalable Feasibility, Impact
LLM-as-a-Judge • Prompt-based judges • Fine-tuned judges • Reinforcement-optimized reward models • Hybrid frameworks LiveIdeaBench (Ruan et al., 2024), CycleReviewer (Weng et al., 2025a), DeepReviewer (Zhu et al., 2025b) HARPA (Vasu et al., 2025), ReviewRL (Zeng et al., 2025) + Scalable, reproducible, approximates expert reasoning and intuition - Sensitive to prompts, training bias, and domain drift Feasibility, Novelty, Impact
Comput-ation Based • Text similarity metrics: BLEU, BERTScore, DSI • Semantic distance: RND, HD, ON, APD, Vendi Score • Graph centrality • Bayesian fusion SciMon (Wang et al., 2024a), CrPO (Ismayilzada et al., 2025a), VirSci (Su et al., 2024), RND (Wang et al., 2025e), Vendi Score (Friedman and Dieng, 2023), SciND (Gupta et al., 2024; Ke et al., 2023), HypoAgent (Duan et al., 2025) + Objective, interpretable, domain-agnostic, scalable - Surface-level semantics; depends on corpus quality and embedding fidelity Novelty, Diversity, Impact
Execution Based • Simulators, compilers, autonomous labs • Behavioral execution AI Scientist (Lu et al., 2024a), AI Co-Scientist (Gottweis et al., 2025), MindShift (Wu et al., 2024), LLM Counter Speech (Podolak et al., 2024) + Grounded, outcome-based validation - Expensive, domain-limited, low scalability Feasibility, Impact
@@ 7.1.1 Novelty and Originality
Several recent studies have approached novelty assessment through semantic local density, assuming that truly novel ideas should lie in sparse regions of the scientific embedding space. Su et al. (2024) introduced two such measures: Historical Dissimilarity (HD)—the average Euclidean distance between a generated idea’s embedding and its five most similar historical abstracts—and Overall Novelty (ON), which further incorporates the contemporaneity and impact of neighboring works:
@@
Building on this intuition, Relative Neighbor Density (RND) evaluates novelty through the density of an idea’s semantic neighborhood relative to its neighbors’ own densities (Wang et al., 2025e). Given a set of ideas $I={{\text{idea}_{i}}}$, each represented by an embedding $v_{i}=G(\text{idea}_{i})$ obtained via a semantic encoder $G(\cdot)$, RND computes for each idea $i$ its $P$ nearest neighbors in the literature corpus $A$:
@@ 7.1.2 Diversity
While novelty measures how far a single idea departs from prior knowledge, diversity evaluates the breadth of a set of generated ideas. A model that outputs only a narrow cluster of novel but similar hypotheses may underexplore the solution space. Diversity thus ensures that systems span multiple conceptual directions, not just isolated rare points. Empirical studies confirm that high novelty does not imply diversity—models may generate unique yet domain-concentrated ideas—underscoring the need to evaluate both dimensions jointly (Ruan et al., 2024).
@@
Building on the same principle, instead of evaluating the whole set at once, CrPO Ismayilzada et al. (2025a), computes a per-response dispersion score. Responses with higher score—i.e., those further from their peers—receive stronger rewards, encouraging models to cover multiple directions rather than collapsing to a single theme.
Beyond cosine-based measures, Vendi Score Friedman and Dieng (2023) provides a more principled measure of set distinctness using Shannon entropy over the eigenvalues of the similarity matrix. Given a similarity matrix $K\in\mathbb{R}^{M\times M}$ constructed from embeddings, let $\{\lambda_{1},\dots,\lambda_{M}\}$ denote its normalized eigenvalues (summing to 1). The Vendi Score is defined as:
@@ 7.1.3 Feasibility and Resource-Efficiency
The HypoAgents framework (Duan et al., 2025) proposes an N–R–F (Novelty–Relevance–Feasibility) model that quantifies idea quality via Bayesian inference and information fusion. Using retrieved literature and semantic embeddings, it measures feasibility as the Bayesian probability of supporting evidence—deriving scores computationally.
Execution-grounded evaluations appear most clearly in autonomous research pipelines. For instance, AI Scientist papers (Lu et al., 2024a; Yamada et al., 2025) or DeepScientist (Weng et al., 2025b) integrates idea generation, code synthesis, and experimental validation in a closed loop. Hypotheses proposed by one agent are automatically implemented as machine-learning experiments, trained, and benchmarked; success metrics such as performance gain or reproducibility indicate feasibility. Similarly, HARPA (Vasu et al., 2025) introduces a testability-driven ideation framework, where an LLM reward model predicts feasibility scores and selected hypotheses are subsequently verified through simulation or real-world trials, closing the loop between idea generation and empirical testing.
In materials and molecular science, execution-based feasibility is operationalized through simulation and experimental pipelines. Borg et al. (2023) introduce empirical measures such as Discovery Yield—the fraction of target materials discovered—and Discovery Probability—the likelihood of identifying at least one viable candidate—both derived from iterative experimental cycles. Similarly, Gallagher and Webb (2025) quantify efficiency as the proportion of candidate compounds passing synthesizability, stability, or toxicity constraints during virtual screening, effectively measuring the fraction of computationally or experimentally “non-wasted” recommendations. Systems like BenchML (Poelking et al., 2021) combine simulators and synthesis predictors to test chemical hypotheses, evaluating feasibility based on success rates such as molecular stability or docking performance.
@@ 7.1.4 Utility and Impact
where $\sigma_{st}$ is the total number of shortest paths between nodes $s$ and $t$, and $\sigma_{st}(y)$ is the number of such paths passing through $y$. High $C_{B}(y)$ indicates that the idea bridges otherwise disconnected communities, a property correlated with transformative scientific discoveries (Ke et al., 2023). Such metrics provide a scalable, domain-agnostic way to estimate potential influence before real deployment, though they cannot capture context-specific or practical impact.
@@
The AI Co-Scientist (Gottweis et al., 2025) explicitly executes biomedical hypotheses to quantify real impact through end-to-end wet-lab experimentation. It autonomously proposes and tests hypotheses in three increasingly complex domains—drug repurposing, novel target discovery, and microbial evolution. The system systematically measures downstream biological outcomes—such as inhibition efficacy or regenerative response—to assess the practical significance of its discoveries. This rigorous execution-driven evaluation exemplifies how impact in scientific ideation can be assessed through direct experimentation rather than proxy metrics.
Beyond laboratory science, behavioral execution studies such as MindShift (Wu et al., 2024) and LLM Counter-Speech (Podolak et al., 2024) deploy AI-generated interventions in natural environments—digital health and social media—to quantify real behavioral change as a domain-specific key performance indicator (KPI) measuring practical impact. Similarly, large-scale agent-based frameworks like AgentSociety (Piao et al., 2025) model how new social or policy ideas propagate through simulated societies, tracking collective outcomes as impact proxies. In summary, impact assessment through execution-driven tests, behavioral outcome studies or graph-based scientometric, ensure that models are not merely producing ideas that are novel, diverse, or feasible, but ones that are also meaningful and capable of advancing science or society in a measurable way. However, such approaches are inherently limited: many scientific ideas are not directly executable, domain simulators exist only for select fields, and measurable outcomes often capture only the practical success of an idea rather than its deeper creative or theoretical significance. Consequently, nuanced dimensions of creativity and long-term scientific value still rely on human expertise and intuition, which remain the gold standard for assessing scientific ideas.
Table 4 summarizes the main evaluation paradigms—from human judgment to computational and execution-based methods—and their roles in assessing novelty, diversity, feasibility and impact.
@@ 7.2 Human Expert Evaluation
Typical setups follow a panel review format, modeled after peer review. For example, AIdeation (Wang et al., 2025d) engaged professional concept designers to rate generated ideas using Likert scales (e.g., 1–5 ratings) for creativity, feasibility, and satisfaction, supplemented by qualitative interviews. Similar setups are found in Scideator (Smith and Huang, 2025) and Nova (Hu et al., 2024), where domain specialists score hypotheses for novelty, plausibility, and usefulness, providing both quantitative and qualitative human feedback. More recently, large-scale human studies have emerged to improve robustness and generalizability of human evaluation. Notably, Si et al. (2024) evaluates LLM-generated research ideas using assessments from over 100 domain experts, explicitly separating judgments of novelty and usefulness.
To improve inter-rater reliability, several studies adopt the Consensual Assessment Technique (CAT) (Amabile, 1982), aggregating multiple expert scores into consensus ratings that reduce bias and align evaluation with established creativity research. At the most rigorous level, AI Scientist v2 (Yamada et al., 2025) elevates human evaluation to formal peer review by submitting autonomously generated papers to top-tier conferences, where acceptance serves as the strongest validation of impact and scientific merit.
@@
Judge category Core idea Example works Strengths Limitations
Inference-time Judges (Prompt-based) Leverages structured prompts, retrieval engines, or multi-agent critique to approximate expert review without fine-tuning. AI Scientist (Lu et al., 2024a), LiveIdeaBench (Ruan et al., 2024), Chain-of-Idea (Li et al., 2024a) Low-cost, interpretable, and reproducible; easy to generalize across domains and tasks. Highly sensitive to prompt phrasing and context; lacks consistent calibration across runs.
Supervised Fine-tuned Judges Trains models on annotated peer-review corpora to internalize expert-like judgment and scoring patterns. CycleReviewer (Weng et al., 2025a), OpenReviewer (Idahl and Ahmadi, 2025) High correlation with expert judgments; improved consistency and interpretability. Memorization instead of reasoning; static preferences; dependent on dataset coverage and quality; may inherit bias.
RL and Reward-Tuned Judges Optimizes evaluative behavior via reinforcement learning or preference alignment for reasoning stability. HARPA (Vasu et al., 2025), ReviewRL (Zeng et al., 2025), REMOR (Taechoyotin and Acuna, 2025) Generalization and adaptivity; high correlation with expert judgments; encourages coherent, stable, and self-consistent evaluations. Expensive to train; vulnerable to reward mis-specification and preference drift.
Hybrid (Training + Inference Integration) Combines fine-tuning with inference-time reasoning and external verifiers or factuality tools. DeepReview (Zhu et al., 2025b), HARPA (Vasu et al., 2025) Merges the generalization of training-based methods with the factual grounding of inference scaling methods. Complex pipeline; requires maintaining multi-stage logic and external dependencies.